OpenStudy (abmon98):

write down the expansion in ascending powers of x up to the term x^3 of i)(1+x)^1/2 ii)(1-x)^-1/2 Hence or otherwise, expand ((1+x)/(1-x))^1/2 in ascending powers of x to the term x^2 by using x=1/10 obtain an estimate to three decimal place for square root of 11

OpenStudy (abmon98):

i)1+x/2-x^2/8+x^3/16 ii)1+x/2+3x^2/8 +5x^3/16

OpenStudy (abmon98):

but how can i estimate the value of square 11 using this expansion

OpenStudy (anonymous):

You're talking about Taylor expansions, right?

OpenStudy (abmon98):

i dont know whats taylors expansion.

OpenStudy (anonymous):

It's a method of expressing a function with continuous derivatives as a series of other functions. For example, you could write \(\sin x\) as an infinite-term polynomial, \[\sin x=x-\frac{x^3}{3!}+\frac{x^5}{5!}-\cdots=\sum_{n=0}^\infty \frac{(-1)^{n}x^{2n+1}}{(2n+1)!}\] Chances are you haven't seen this unless you've taken a course in calculus. Perhaps you know about the binomial theorem?

OpenStudy (abmon98):

yes i do know about binomial theorm

OpenStudy (anonymous):

Okay, in that case, you must be using the expansion formula, \[(1+x)^n=1+nx+\frac{n(n-1)}{2!}x^2+\frac{n(n-1)(n-2)}{3!}x^3+\cdots\] for \(-1<x<1\).

OpenStudy (abmon98):

i did solve part i and part ii and i wrote both answers at the top using this method but iam stuck on the estimation part

OpenStudy (anonymous):

Right, and the expansion of \(\left(\dfrac{1+x}{1-x}\right)^{1/2}\) shouldn't be hard to find. Do you have that too?

OpenStudy (abmon98):

i know that would expand (1+x) and (1-x) seperately but (x-1)^1/2 they are both similar to the expansion of i and ii

OpenStudy (abmon98):

1+x/2-x^2/8/1-x/2-x^2/8

OpenStudy (abmon98):

Substitute x=1/10 and i got 0.9975

OpenStudy (anonymous):

You'll need the expansion for the estimate part. Notice that for \(x=\dfrac{1}{10}\), you have \[\left(\dfrac{1+\dfrac{1}{10}}{1-\dfrac{1}{10}}\right)^{1/2}=\left(\dfrac{\dfrac{11}{10}}{\dfrac{9}{10}}\right)^{1/2}=\left(\frac{11}{9}\right)^{1/2}=\frac{1}{3}\sqrt{11}\] This means that when you plug in \(x=\dfrac{1}{10}\) into the expansion, you have \[\frac{1}{3}\sqrt{11}\approx 1+(\cdots)\frac{1}{10}+(\cdots)\frac{1}{10^2}\] which means \[\sqrt{11}\approx 3\left(1+(\cdots)\frac{1}{10}+(\cdots)\frac{1}{10^2}\right)\]

OpenStudy (abmon98):

why dont we plug x=1/10 directly in the expansion and then do the square root and dividing.

OpenStudy (anonymous):

I don't think your third expansion is right. The first two, \[(1+x)^{1/2}\approx 1+\frac{x}{2}-\frac{x^2}{8}+\frac{x^3}{16}\\ (1-x)^{-1/2}\approx 1+\frac{x}{2}+\frac{3x^2}{8}+\frac{5x^3}{16}\] are correct. The third expansion is the product of the two above: \[\left(\frac{1+x}{1-x}\right)^{1/2}=(1+x)^{1/2}(1-x)^{-1/2}\] so \[\left(\frac{1+x}{1-x}\right)^{1/2}\approx \left(1+\frac{x}{2}-\frac{x^2}{8}+\frac{x^3}{16}\right)\left(1+\frac{x}{2}+\frac{3x^2}{8}+\frac{5x^3}{16}\right)\] Up to the \(x^2\) term, you can ignore the \(x^3\) terms of the component expansions: \[\begin{align*}\left(\frac{1+x}{1-x}\right)^{1/2}&\approx \left(1+\frac{x}{2}-\frac{x^2}{8}\right)\left(1+\frac{x}{2}+\frac{3x^2}{8}\right)\\ &\approx 1+\frac{x}{2}+\frac{3x^2}{8}+\frac{x}{2}+\frac{x^2}{4}+\frac{3x^3}{16}-\frac{x^2}{8}-\frac{x^3}{16}-\frac{3x^4}{64}\\ &\approx 1+\frac{x}{2}+\frac{3x^2}{8}+\frac{x}{2}+\frac{x^2}{4}-\frac{x^2}{8}\\ &\approx 1+x+\frac{x^2}{2} \end{align*}\] So, \[\begin{align*}\frac{1}{3}\sqrt{11}&\approx 1+\frac{1}{10}+\frac{1}{200}\\ \sqrt{11}&\approx 3.315 \end{align*}\]

OpenStudy (abmon98):

Thank you so much :)

OpenStudy (anonymous):

You're welcome, I hope it helped!

OpenStudy (abmon98):

yes it did help but one more question iam sorry how did you get 1/3(11)^1/2

OpenStudy (anonymous):

That's the value you get for plugging \(x=\dfrac{1}{10}\) in the product expression, \((1+x)^{1/2}(1-x)^{-1/2}\).

OpenStudy (abmon98):

oh, thank you again :)

OpenStudy (anonymous):

You're welcome